(a) Rowanberries contain 6,2% sugar. Nine-tenths of the rest of the material is water. The container contains 500 g of rowanberries. How many grams of water do the berries in the container contain? 


b) Up to 32.7% of the fox's hair may be on the fox's tail. If the fox's tail has a total of 15,400 hairs, how many hairs are there elsewhere in the fox's turkey? 


c) A small entrepreneur sells the jewelry he or she makes at a price of 12 euros per piece, which means that the average sales are 340 pieces per month. He will lower the price of jewelery by 11%, increasing his monthly jewelery sales by 5%. What percentage of an entrepreneur’s income will change? 


1
Expert's answer
2021-12-20T13:38:03-0500

(a)


"(1-0.062)\cdot0.9\cdot500\ g=422.1\ g"

The berries in the container contain 422.1 grams of water.


(b)


"\dfrac{100-32.7}{32.7}\cdot15400\approx31695"

(c)


"12(340)=4080 (euros)"

"12(1-0.11)(340)(1+0.05)=3812.76(euros)"

"100\%(\dfrac{4080-3812.76}{4080})=6.55\%"

Entrepreneur’s income will decrease by "6.55\%."